# A Glitch in the Matrix? Locating and Detecting Language Model Grounding with Fakepedia

### v2

@@ Abstract
Large language models (LLMs) have an impressive ability to draw on novel information supplied in their context. Yet the mechanisms underlying this contextual grounding remain unknown, especially in situations where contextual information contradicts factual knowledge stored in the parameters, which LLMs also excel at recalling. Favoring the contextual information is critical for retrieval-augmented generation methods, which enrich the context with up-to-date information, hoping that grounding can rectify outdated or noisy stored knowledge. We present a novel method to study grounding abilities using Fakepedia, a dataset of counterfactual texts constructed to clash with a model’s internal parametric knowledge. We benchmark various LLMs with Fakepedia and then we conduct a causal mediation analysis, based on our Masked Grouped Causal Tracing (MGCT), on LLM components when answering Fakepedia queries. Within this analysis, we identify distinct computational patterns between grounded and ungrounded responses. We finally demonstrate that distinguishing grounded from ungrounded responses is achievable through computational analysis alone. Our results, together with existing findings about factual recall mechanisms, provide a coherent narrative of how grounding and factual recall mechanisms interact within LLMs.
@@ 2.1 Mechanistic Interpretability
Discovering the low-level mechanisms that give rise to high-level behaviors has become an important goal for research in AI interpretability and allowed for better understanding and prediction of models behavior Mu and Andreas (2020); Geiger et al. (2022b); Wu et al. (2023b); Vig et al. (2020); Dai et al. (2022); Oba et al. (2021); Peyrard et al. (2021).
@@
While recalling factual knowledge relates to memorization, in this work we study the ability of models to ground their answers based on information in the context, i.e., to incorporate new information not seen at training time. In particular, we craft scenarios that set the LLMs in tension between the mode of factual recall and the mode of grounding. This contributes to the broader discussion on generalization versus memorization Razeghi et al. (2022); Kandpal et al. (2023); Hupkes et al. (2023); Haviv et al. (2023). However, grounding remains much less studied than factual recall. In a contemporary study, Yu et al. (2023) also inspect the problem of grounding using mechanistic interpretability methods. Their analysis focuses on the role of attention heads when forcing the model to ground its answer in the prompt. Our findings nicely complement theirs. When combined with existing findings about factual knowledge and information flow during recall, our results begin to portray a coherent narrative that we detail in Sec. 8.
@@ 6.1 MGCT Analysis
Figure 2: Masked Grouped Causal Tracing (MGCT). This figure illustrates the mediation analysis from MGCT, refining the preceding causal tracing method from Meng et al. (2023a). The process involves three steps: (i) Clean run: all states within the computation graph are recorded during a forward loop, resulting in a predicted token, in this case “Paris.” (ii) Corrupted run: the subject tokens are substituted with special non-textual tokens such as <UNK> or <EOS>, leading to a distinct probability for the predicted token. (iii) Restored run: the corrupted with the restoration of a group of states (in this instance, four hidden activations) to their values from the clean run, resulting in a partially restored probability for the predicted token. The indirect effect is estimated by the extent to which the restoration of these states contributes to the probability restoration of the predicted token.
@@
Masked Causal Tracing. In our experiments, we observed that the outcome of the causal tracing algorithm is highly sensitive to the choice and magnitude of noise in corrupted runs. To enhance the robustness and generalizability of causal tracing, we propose to perturb directly the tokens instead of the embeddings. This corruption method involves substituting tokens with special non-textual tokens, such as the UNK or EOS tokens. This modification results in more stable masked causal traces.
@@ 6.2 MGCT Experiments
Experimental setup. We perform MGCT analysis on GPT2-XL (Radford et al., 2019) and Llama2-7B (Touvron et al., 2023) using the Fakepedia dataset. Data points were categorized as grounded or ungrounded based on model answers, considering only the first token for instances with multiple tokens. The analysis involved simultaneous restoration of all states across all layers for a specific token, reducing the computational time by orders of magnitude compared to Meng et al. (2023a).
@@ 7 Automatic Detection of Ungrounded Responses
Within our MGCT analysis, we identify distinct computation patterns between grounded and ungrounded responses. To automate detection, we create a balanced dataset with 4,000 GPT2-XL responses from the Fakepedia-Base dataset. Employing an 80%-20% training-test split, we extract 18 features from MGCT outputs for a binary classifier (e.g., attention, MLP, hidden activation effects, by grouping tokens results as in the MGCT mediation analysis). After training an XGBoost classifier, we achieve a 92.8% accuracy on the test set. Feature importance analysis identifies MLPs on the last subject tokens as most crucial, aligning with our MGCT findings. Our study demonstrates that distinguishing grounded from ungrounded responses is achievable through computation analysis alone. We refer to Appendix E for a more detailed description of the experimental setup.
@@ B.2 Fakepedia
[…]

### v3

@@ Abstract
Large language models (LLMs) have an impressive ability to draw on novel information supplied in their context. Yet the mechanisms underlying this contextual grounding remain unknown, especially in situations where contextual information contradicts factual knowledge stored in the parameters, which LLMs also excel at recalling. Favoring the contextual information is critical for retrieval-augmented generation methods, which enrich the context with up-to-date information, hoping that grounding can rectify outdated or noisy stored knowledge. We present a novel method to study grounding abilities using Fakepedia, a novel dataset of counterfactual texts constructed to clash with a model’s internal parametric knowledge. We benchmark various LLMs with Fakepedia and conduct a causal mediation analysis of LLM components when answering Fakepedia queries, based on our Masked Grouped Causal Tracing (MGCT) method. Through this analysis, we identify distinct computational patterns between grounded and ungrounded responses. We finally demonstrate that distinguishing grounded from ungrounded responses is achievable through computational analysis alone. Our results, together with existing findings about factual recall mechanisms, provide a coherent narrative of how grounding and factual recall mechanisms interact within LLMs.
[…]
FakepediaExample O!htbp m floatplacement=#1, float, width=colback=gray!5!white, colframe=gray!75!black, title=#2,
A Glitch in the Matrix?
Locating and Detecting Language Model Grounding with Fakepedia
Giovanni Monea, Maxime Peyrard, Martin Josifoski, Vishrav Chaudhary, Jason Eisner, Emre Kıcıman, Hamid Palangi, Barun Patra, Robert West EPFL  Univ. Grenoble Alpes, CNRS, Grenoble INP, LIG  Microsoft Corporation {giovanni.monea, martin.josifoski, robert.west}@epfl.ch {maxime.peyrard}@univ-grenoble-alpes.fr {vchaudhary, jason.eisner, emrek, hpalangi, barun.patra}@microsoft.com
@@ 2.1 Mechanistic Interpretability
Discovering the low-level mechanisms that give rise to high-level behaviors has become an important goal for research in AI interpretability and allowed for better understanding and prediction of models behavior Mu and Andreas (2020); Geiger et al. (2022b); Wu et al. (2023b); Vig et al. (2020); Dai et al. (2022); Oba et al. (2021); Peyrard et al. (2021). Understanding these mechanisms allows us to better predict out-of-distribution behavior Mu and Andreas (2020); Geiger et al. (2022b); Wu et al. (2023b), identify and fix errors Vig et al. (2020). For instance, Dai et al. (2022) analyzed BERT activations revealing that some neurons are positively correlated with specific facts. Similarly, Oba et al. (2021) demonstrated associations between neuron activations and specific phrases in model output. Additionally, in a matched study on humor detection, Peyrard et al. (2021) identified attention heads in BERT encoding the funniness of sentences.
@@
While recalling factual knowledge relates to memorization, in this work we study the ability of models to ground their answers based on information in the context, i.e., to incorporate new information not seen at training time. In particular, we craft scenarios that set the LLMs in tension between the mode of factual recall and the mode of grounding. This contributes to the broader discussion on generalization versus memorization Razeghi et al. (2022); Kandpal et al. (2023); Hupkes et al. (2023); Haviv et al. (2023); Xu et al. (2024). However, grounding remains much less studied than factual recall. In a contemporary study, Yu et al. (2023) also inspect the problem of grounding using mechanistic interpretability methods. Their analysis focuses on the role of attention heads when forcing the model to ground its answer in the prompt. Our findings nicely complement theirs. When combined with existing findings about factual knowledge and information flow during recall, our results begin to portray a coherent narrative that we detail in Sec. 8.
@@ 6.1 MGCT Analysis
Figure 2: Masked Grouped Causal Tracing (MGCT). This figure illustrates the mediation analysis from MGCT, refining the preceding causal tracing method from Meng et al. (2023a). The process involves three steps: (i) Clean run: all states within the computation graph are recorded during a forward loop, resulting in a predicted token, in this case “Paris”. (ii) Corrupted run: the subject tokens are substituted with special non-textual tokens such as <UNK> or <EOS>, leading to a distinct probability for the predicted token. (iii) Restored run: the corrupted with the restoration of a group of states (in this instance, four hidden activations) to their values from the clean run, resulting in a partially restored probability for the predicted token. The indirect effect is estimated by the extent to which the restoration of these states contributes to the probability restoration of the predicted token.
@@
Masked causal tracing. In our experiments, we observed that the outcome of the causal tracing algorithm is highly sensitive to the choice and magnitude of noise in corrupted runs. To enhance the robustness and generalizability of causal tracing, we propose to perturb directly the tokens instead of the embeddings. This corruption method involves substituting tokens with special non-textual tokens, such as the UNK or EOS tokens. This modification results in more stable masked causal traces, avoids the need for multiple runs to average out the randomness, and consequently requires less memory and increases the speed of the process.
@@ 6.2 MGCT Experiments
Experimental setup. We perform MGCT analysis on GPT2-XL (Radford et al., 2019) and Llama2-7B (Touvron et al., 2023) using the Fakepedia dataset. Data points were categorized as grounded or ungrounded based on model answers, considering only the first token for instances with multiple tokens. The analysis involved simultaneous restoration of all states across all layers for a specific token, reducing the computational time by orders of magnitude compared to Meng et al. (2023a). This approach effectively examines the impact of each token on the output.
@@ 7 Automatic Detection of Ungrounded Responses
Within our MGCT analysis, we identify distinct computation patterns between grounded and ungrounded responses. To automate detection, we create a balanced dataset with 4,000 GPT2-XL responses from the Fakepedia-base dataset. Employing an 80%–20% train–test split, we extract 18 features from MGCT outputs for a binary classifier (e.g., attention, MLP, hidden activation effects, by grouping tokens results as in the MGCT mediation analysis). After training an XGBoost classifier, we achieve a 92.8% accuracy on the test set. Feature importance analysis identifies MLPs on the last subject tokens as most crucial, aligning with our MGCT findings. Our study demonstrates that distinguishing grounded from ungrounded responses is achievable through computation analysis alone. We refer to Appendix E for a more detailed description of the experimental setup.
@@ B.1 Counterfactual ParaRel
Table 2: Accuracy of various LLMs in identifying true objects from Fakepedia facts.
Mistral Zephyr Llama2 GPT-3.5 Turbo GPT-4 Turbo
Dataset 7B 7B 7B 13B 70B 03/01 06/13 11/06 11/06
FP 99.4% 99.5% 94.4% 97.5% 98.9% 99.9% 99.9% 99.9% 100%
FP-MH 98.4% 98.8% 91.9% 98.4% 96.2% 99.6% 99.6% 99.5% 99.9%
Do all models know about the facts selected by GPT2?
We also verify that all the other models used in our analyses similarly know and can recall the facts selected by GPT2-XL, as previously outlined. In particular, for the models used in Sec. 5, we maintain the same experimental setup while modifying the prompt to exclude the context and make the models choose between the true and false objects without any context. The results are summarized in Table 2. All the models can choose the true objects in more than 90% of the cases, although this figure increases to more than 98% for most models. Likewise, we verify that Llama2-7B, used in Sec. 6, can correctly assign high probabilities to the true objects as the next tokens, in the same way GPT2-XL did when choosing the facts in the first place. We verified that the true objects are within the three most likely next tokens in 99.4% of the cases (95% for the most likely, 3.7% for the second most likely, and 0.7% for the third most likely) on Fakepedia-base. The respective figures for Fakepedia-MH are 98.8% (90.4%, 6.9%, 1.5%).
@@ B.2 Fakepedia
Fakepedia-MH. In the multi-hop variant, our objective is to produce textual descriptions that do not explicitly state the triplet but logically imply it. This approach tests the model’s ability not only to extract information from context but also to integrate and engage in basic reasoning to derive the answer. When composing a 2-hop paragraph description for the triplet (subja, rela, obja), we rely on an intermediary triplet from Fakepedia-base. Given a triplet (subjb, rela, obja) from Fakepedia-base, we select at random the target triplet (subja, rela, obja) from the counterfactual triplets with the same relation (rela) and the same object (obja). Given these two triplets, we use the Fakepedia-base description for the selected intermediary triplet (subjb, rela, obja), and append a linking senetence that logically implies the target triplet (subja, rela, obja). The linking sentence is generated from a template, such as [X] and [Y] belong to the same continent, that we adapt from ParaRel templates for queries. We publicly release all the new templates with our code.